# section[1.5em]1.5em1pc

### v2

@@ 7.   Derivations, Module of Kähler differentials
Let $M$ be an $A$-module, and ${B\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{pr}$${A}$ be a $\Gamma(\mathcal{P})$-algebra over $A$. A Beck derivation, or simply derivation from $B$ to $M$ is a linear map $d\colon B\to M$ such that $pr+d\colon B\to A\ltimes M$ is a morphism of $\Gamma(\mathcal{P})$-algebra. We denote by $\operatorname{Der}_{A}(B,M)$ the vector space of derivations from $B$ to $M$. We obtain a bifunctor
@@
A Beck derivation $d\colon B\to M$ is a linear map such that:
@@
$(pr+d)(\beta_{x,{\underline{r}}}\left(b_{1},\dots,b_{s}\right))=\beta_{x,{\underline{r}}}\left((pr(b_{1}),d(b_{1})),\dots,(pr(b_{s}),d(b_{s}))\right)\\ =\left(\beta_{x,{\underline{r}}}\left(pr(b_{1}),\dots,pr(b_{s})\right),\sum_{j=1}^{s}\sum_{l,l^{\prime}}\beta_{x,{\underline{r}}\circ_{j}(l,l^{\prime})}(pr(b_{1}),\dots,pr(b_{j}),d(b_{j}),pr(b_{j+1}),\dots,pr(b_{s})).\right),$
@@
$\beta_{x,{\underline{r}}}(a_{1},\dots,a_{s-1},-)\otimes\lambda\beta_{y,{\underline{q}}}(b_{1},\dots,b_{u})\mapsto\\ \lambda^{r_{s}}\sum_{j=1}^{u}\sum_{l+l^{\prime}=q_{j},l^{\prime}>0}\mu\left(\beta_{x,{\underline{r}}}(a_{1},\dots,a_{s-1},-)\otimes\beta_{y,{\underline{q}}\circ_{j}(l,l^{\prime})}(b_{1},\dots,b_{j},-,b_{j+1},\dots,b_{u})\right)\otimes db_{j},$
@@
Let us describe $\Omega_{\Gamma(\mathcal{P})}(A)$ in more detail. For an element of the type $\beta_{x,{\underline{r}}}(a_{1},\dots,a_{a-1},-)\otimes da\in\mathbb{U}_{\Gamma(\mathcal{P})}(A)\otimes dA$, we will denote by $\beta_{x,{\underline{r}}}(a_{1},\dots,a_{a-1},da)$ its image in $\Omega_{\Gamma(\mathcal{P})}(A)$. Following 4.6, we will allow the notation:
$\beta_{x,{\underline{r}}}(a_{1},\dots,a_{i-1},da,a_{i+1},\dots,a_{s}):=\beta_{\sigma_{i}^{*},{\underline{r}}^{\sigma_{i}}}(a_{1},\dots,a_{i-1},a_{i+1},\dots,a_{s},da).$
@@
$\Omega_{\Gamma(\mathcal{P})}(A)$ represents the functor $\operatorname{Der}(A,-)\colon A\text{-}\mathrm{Mod}\to\operatorname{Ab}$ sending $M$ to the abelian group of derivations from $A$ to $M$.
@@
We have to show that for all $A$-module $M$, there is a linear bijection, natural in $M$:
@@
Let $f\colon\Omega_{\Gamma(\mathcal{P})}(A)\to M$ be an $A$-module morphism. Consider the linear map $D\colon A\to M$ given by $D=f\circ d$, where $d\colon A\to\Omega_{\Gamma(\mathcal{P})}(A)$ is the universal derivation. Since the universal derivation is a derivation, and since $f$ is an $A$-module morphism, $D$ is a derivation. The assignment $f\mapsto D$ yields a linear map $\phi\colon\operatorname{Hom}_{A\text{-}\mathrm{Mod}}(\Omega_{\Gamma(\mathcal{P})}(A),M)\to\operatorname{Der}_{A}(A,M)$, natural in $M$.
Conversely, let $D\colon A\to M$ be a derivation. We can consider $D$ as a linear map $dA\to M$. This extends uniquely into an $A$-module morphism $\mathbb{U}_{\Gamma(\mathcal{P})}(A)\otimes dA\to M$. The fact that $D$ is a derivation ensures that this passes to the coequalizer into an $A$-module morphism $f\colon\Omega_{\Gamma(\mathcal{P})}(A)\to M$. This assignment $D\mapsto f$ yields a linear map $\psi\colon\operatorname{Der}_{A}(A,M)\to\operatorname{Hom}_{A\text{-}\mathrm{Mod}}(\Omega_{\Gamma(\mathcal{P})}(A),M)$. It is straightforward to check that $\phi$ and $\psi$ are mutually inverse and natural in $M$. ∎
To conclude this section, let us link this new notion of Kähler differentials for a $\Gamma(\mathcal{P})$-algebra to the usual notion of Kähler differentials for a $\mathcal{P}$-algebra. Recall that $U^{\Gamma(\mathcal{P})}_{\mathcal{P}}\colon\Gamma(\mathcal{P})\text{-}\mathrm{Alg}\to\mathcal{P}\text{-}\mathrm{Alg}$ denotes the forgetful functor from $\Gamma(\mathcal{P})$-algebras to $\mathcal{P}$-algebras. Then, if $A$ is a $\Gamma(\mathcal{P})$-algebra, $\Omega_{\mathcal{P}}(A):=\Omega_{\mathcal{P}}(U^{\Gamma(\mathcal{P})}_{\mathcal{P}}(A))$ is the usual $U^{\Gamma(\mathcal{P})}_{\mathcal{P}}(A)$-module of Kähler differentials of the $\mathcal{P}$-algebra $U^{\Gamma(\mathcal{P})}_{\mathcal{P}}(A)$, as in [LodayV12]*12.3.8.
@@ 8.   Quillen cohomology
Let $f\colon B\to A$ be a morphism of $\Gamma(\mathcal{P})$-algebras. Via the equivalence from 5.2, the pushforward along $f$ is given by $f_{!}(M)=\mathbb{U}_{\Gamma(\mathcal{P})}(A)\otimes_{\mathbb{U}_{\Gamma(\mathcal{P})}(B)}M$.
@@
The universal enveloping algebra $\mathbb{U}_{\Gamma(\mathcal{P})}$ provides a functor $\Gamma(\mathcal{P})\text{-}\mathrm{Alg}\to\operatorname{As}_{+}\text{-}\mathrm{Alg}$ to unital rings. The ring homomorphism $\mathbb{U}_{\Gamma(\mathcal{P})}(f)\colon\mathbb{U}_{\Gamma(\mathcal{P})}(B)\to\mathbb{U}_{\Gamma(\mathcal{P})}(A)$ makes $\mathbb{U}_{\Gamma(\mathcal{P})}(A)$ into a right $\mathbb{U}_{\Gamma(\mathcal{P})}(B)$-module. Via the equivalence from 6.11, the pushforward adjunction
@@
(1) Via the equivalences from 5.2 and 6.11, the statement follows from combining Proposition 7.6, 8.1, and 2.4.
(2) Via the equivalence from 6.11, the right adjoint $A\ltimes-$ is the forgetful functor
@@
The claim then follows from [Frankland15]*Proposition 3.40. ∎
@@ 9.   Example: The operad $\mathrm{Com}$
In this section, we apply our general construction for Beck modules, universal enveloping algebra, Beck derivations and Kähler differentials for divided power algebras over the operad $\mathrm{Com}$ of associative, commutative (non-unital) algebras. We check that these correspond to the construction obtained by the first author in [Dokas09, Dokas23].
@@ 9.1.   Classical definition and state of the art
Given an augmented $\mathbb{F}$-algebra $A$ with augmentation $\epsilon\colon A\to\mathbb{F}$, denote the augmentation kernel by $A_{+}=\ker(A\to\mathbb{F})$.
@@
We follow the notation $A_{+}=\ker(A\to\mathbb{F})$ used in [Soublin87] and [Dokas09], though some authors use a different notation, notably $\overline{A}$ in [LodayV12]*§1.1.
@@
We define the ring $V(A)$ as the ring whose underlying $\mathbb{F}$-vector space is the tensor product $V(A):=A\otimes_{\mathbb{F}}R_{f}$ and the multiplication is given by:
@@ 9.2.   Operadic point of view
Let us now use the characterisation of $\Gamma(\mathrm{Com})$-algebra from 3.3. Following [Ikonicoff20], if $B$ is a $\Gamma(\mathrm{Com})$-algebra, its multiplication is given by $\beta_{X_{2},(1,1)}$, and $\pi$, which represents the divided power operation $\gamma_{p}$, is given by $\beta_{X_{p},(p)}$.
@@
Note that $B$ injects into $\mathbb{U}_{\Gamma(\mathrm{Com})}(B)$, by identifying $b\in B$ to $\beta_{X_{2},(1,1)}(b,-)$. For a $B$-module (i.e., a $\mathbb{U}_{\Gamma(\mathrm{Com})}(B)$-module) $M$, we will denote by $bm:=\beta_{X_{2},(1,1)}(b,m)=\beta_{X_{2},(1,1)}(b,-)\cdot m$, and $\pi(m):=\beta_{X_{p},(p)}(m)$, for $b\in B$, $m\in M$.
@@
Following Proposition 7.2, a Beck derivation $d\colon B\to M$ is a linear map $d\colon B\to M$ satisfying $d(ab)=adb+bda$ and
@@
Finally, since in characteristic $p$, $(p-1)!=-1$ (see Wilson’s Theorem [DummitF04]*§13.5 Exercise 6), one has:
@@
Following Definition 7.3, the module of Kähler differentials $\Omega_{\Gamma(\mathrm{Com})}(B)$ of $B$ is the $\mathbb{U}_{\Gamma(\mathrm{Com})}(B)$-module generated by elements $db$ for $b\in B$, linear in $b$, under the relations:
$d(ab)$ $=d(\beta_{X_{2},(1,1)}(a,b))=\beta_{X_{2},(1,1)}(-,b)\otimes da+\beta_{X_{2},(1,1)}(a,-)\otimes db=bda+adb,$
@@
$\beta_{X_{p},(p-1,1)}(b,-)\otimes b=\beta_{X_{2},(1,1)}(\beta_{X_{p-1},(p-1)}(b),-)\otimes db=\frac{1}{(p-1)!}b^{p-1}db=-b^{p-1}db.$
@@
For all $i\in\{2,\dots,p-1\}$, we have again $X_{p}=i\frac{1}{i}X_{p}=\sum_{\sigma\in\Sigma_{p-i}\times\Sigma_{i}/\Sigma_{p-i}\times\Sigma_{1}\times\Sigma_{i-1}}\frac{1}{i}X_{p}$. So, $\beta_{X_{p},(p-i,i)}(b,-)=\sum_{\sigma\in\Sigma_{p-i}\times\Sigma_{i}/\Sigma_{p-i}\times\Sigma_{1}\times\Sigma_{i-1}}\frac{1}{i}X_{p}$ is one of the elements described in Definition 6.1 in point (7), and so, is equal to 0 in $\mathbb{U}_{\Gamma(\mathrm{Com})}(B)$. Finally, this last relation reads:
@@ 10.1.   Classical definition and state of the art
Following [Dokas04], denote by $w(L)$ the $\mathbb{F}$-vector space $R_{f}\otimes_{\mathbb{F}}u(L)$, where $R_{f}$ is the polynomial ring on one indeterminate $f$ as in Section 9. Then, $w(L)$ equipped with a ring structure such that $R_{f}\to w(L)$ and $u(L)\to w(L)$ are ring homomorphisms, and such that:
@@
$\operatorname{Der}_{p}(L^{\prime},(M,f)):=\{d\in\operatorname{Der}(L^{\prime},M):d(l^{[p]})=\underbrace{l\cdots l}_{p-1}dl+f(d(l)),\;l\in L^{\prime}\}$
@@
What we call the module of Kähler differentials of the Lie algebra $L$ is, by analogy with the case of commutative algebras, the $L$-Beck module $\Omega_{\operatorname{RLie}}(L)$ which represents the functor of Beck derivations. It follows from 10.8, Proposition 10.10 and [Pareigis68]*Lemma 2.1 that the module $\Omega_{\operatorname{RLie}}(L)$ of Kähler differentials is nothing else than the $w(L)$-module $C(L)$ considered by Pareigis in [Pareigis68], cf. [Dokas04]*§1.3.
@@
The module of Kähler differentials $\Omega_{\operatorname{RLie}}(L)$ for the restricted Lie algebra $L$ is the $w(L)$-module with the following presentation: the generators are the symbols $dl$ for $l\in L$, and the relations are:

### v3

@@ 7.   Derivations, Module of Kähler differentials
Let $M$ be an $A$-module, and ${B\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{pr}$${A}$ be a $\Gamma(\mathcal{P})$-algebra over $A$. A Beck derivation, or simply derivation from $B$ to $M$ is a linear map $d\colon B\to M$ such that $pr+d\colon B\to A\ltimes M$ is a morphism of $\Gamma(\mathcal{P})$-algebras. We denote by $\operatorname{Der}_{A}(B,M)$ the vector space of derivations from $B$ to $M$. We obtain a bifunctor
@@
A Beck derivation is a linear map $d\colon B\to M$ such that:
@@
$(pr+d)(\beta_{x,{\underline{r}}}\left(b_{1},\dots,b_{s}\right))=\beta_{x,{\underline{r}}}\left((pr(b_{1}),d(b_{1})),\dots,(pr(b_{s}),d(b_{s}))\right)$
$=\Bigg{(}\beta_{x,{\underline{r}}}\left(pr(b_{1}),\dots,pr(b_{s})\right),\sum_{j=1}^{s}\sum_{l,l^{\prime}}\beta_{x,{\underline{r}}\circ_{j}(l,l^{\prime})}(pr(b_{1}),\dots,pr(b_{j}),d(b_{j}),pr(b_{j+1}),\dots,pr(b_{s}))\Bigg{)},$
@@
$\beta_{x,{\underline{r}}}(a_{1},\dots,a_{s-1},-)\otimes\lambda\beta_{y,{\underline{q}}}(b_{1},\dots,b_{u})\mapsto$ $\lambda^{r_{s}}\sum_{j=1}^{u}\sum_{l+l^{\prime}=q_{j},l^{\prime}>0}\mu\left(\beta_{x,{\underline{r}}}(a_{1},\dots,a_{s-1},-)\otimes\beta_{y,{\underline{q}}\circ_{j}(l,l^{\prime})}(b_{1},\dots,b_{j},-,b_{j+1},\dots,b_{u})\right)\otimes db_{j},$
@@
Let us describe $\Omega_{\Gamma(\mathcal{P})}(A)$ in more detail. For an element of the type
$\beta_{x,{\underline{r}}}(a_{1},\dots,a_{a-1},-)\otimes da\in\mathbb{U}_{\Gamma(\mathcal{P})}(A)\otimes dA,$
we will denote by $\beta_{x,{\underline{r}}}(a_{1},\dots,a_{a-1},da)$ its image in $\Omega_{\Gamma(\mathcal{P})}(A)$. Following Notation 4.6, we will use the notation:
$\beta_{x,{\underline{r}}}(a_{1},\dots,a_{i-1},da,a_{i+1},\dots,a_{s})\coloneqq\beta_{\sigma_{i}^{*},{\underline{r}}^{\sigma_{i}}}(a_{1},\dots,a_{i-1},a_{i+1},\dots,a_{s},da).$
@@
$\Omega_{\Gamma(\mathcal{P})}(A)$ represents the functor $A\text{-}\mathrm{Mod}\to\operatorname{Ab}$ sending $M$ to the abelian group of derivations from $A$ to $M$.
@@
We have to show that for all $A$-modules $M$, there is a linear bijection, natural in $M$:
@@
Let $f\colon\Omega_{\Gamma(\mathcal{P})}(A)\to M$ be an $A$-module morphism. Consider the linear map $D\colon A\to M$ given by $D=f\circ d$, where $d\colon A\to\Omega_{\Gamma(\mathcal{P})}(A)$ is the universal derivation. Since the universal derivation is a derivation, and since $f$ is an $A$-module morphism, $D$ is a derivation. The assignment $f\mapsto D$ yields a linear map
$\phi\colon\operatorname{Hom}_{A\text{-}\mathrm{Mod}}(\Omega_{\Gamma(\mathcal{P})}(A),M)\to\operatorname{Der}_{A}(A,M),$
natural in $M$.
Conversely, let $D\colon A\to M$ be a derivation. We can consider $D$ as a linear map $dA\to M$. This extends uniquely into an $A$-module morphism $\mathbb{U}_{\Gamma(\mathcal{P})}(A)\otimes dA\to M$. The fact that $D$ is a derivation ensures that this passes to the coequalizer into an $A$-module morphism $f\colon\Omega_{\Gamma(\mathcal{P})}(A)\to M$. This assignment $D\mapsto f$ yields a linear map
$\psi\colon\operatorname{Der}_{A}(A,M)\to\operatorname{Hom}_{A\text{-}\mathrm{Mod}}(\Omega_{\Gamma(\mathcal{P})}(A),M).$
It is straightforward to check that $\phi$ and $\psi$ are mutually inverse and natural in $M$. ∎
To conclude this section, let us link this new notion of Kähler differentials for a $\Gamma(\mathcal{P})$-algebra to the usual notion of Kähler differentials for a $\mathcal{P}$-algebra. Recall that $U^{\Gamma(\mathcal{P})}_{\mathcal{P}}\colon\Gamma(\mathcal{P})\text{-}\mathrm{Alg}\to\mathcal{P}\text{-}\mathrm{Alg}$ denotes the forgetful functor from $\Gamma(\mathcal{P})$-algebras to $\mathcal{P}$-algebras. Then, if $A$ is a $\Gamma(\mathcal{P})$-algebra, $\Omega_{\mathcal{P}}(A)\coloneqq\Omega_{\mathcal{P}}(U^{\Gamma(\mathcal{P})}_{\mathcal{P}}(A))$ is the usual $U^{\Gamma(\mathcal{P})}_{\mathcal{P}}(A)$-module of Kähler differentials of the $\mathcal{P}$-algebra $U^{\Gamma(\mathcal{P})}_{\mathcal{P}}(A)$, as in [LodayV12]*12.3.8.
@@ 8.   Quillen cohomology
Let $f\colon B\to A$ be a morphism of $\Gamma(\mathcal{P})$-algebras. Via the equivalence from Theorem 5.2, the pushforward along $f$ is given by $f_{!}(M)=\mathbb{U}_{\Gamma(\mathcal{P})}(A)\otimes_{\mathbb{U}_{\Gamma(\mathcal{P})}(B)}M$.
@@
The universal enveloping algebra provides a functor
$\mathbb{U}_{\Gamma(\mathcal{P})}\colon\Gamma(\mathcal{P})\text{-}\mathrm{Alg}\to\operatorname{As}_{+}\text{-}\mathrm{Alg}$
to unital rings. The ring homomorphism $\mathbb{U}_{\Gamma(\mathcal{P})}(f)\colon\mathbb{U}_{\Gamma(\mathcal{P})}(B)\to\mathbb{U}_{\Gamma(\mathcal{P})}(A)$ makes $\mathbb{U}_{\Gamma(\mathcal{P})}(A)$ into a right $\mathbb{U}_{\Gamma(\mathcal{P})}(B)$-module. Via the equivalence from Theorem 6.11, the pushforward adjunction
@@
  3. (3)
    Let $C_{\bullet}\xrightarrow{\sim}A$ be a simplicial resolution of $A$, i.e., a cofibrant replacement in $s(\Gamma(\mathcal{P})\text{-}\mathrm{Alg})$. The cotangent complex of $A$ is the simplicial $A$-module
    $\mathbf{L}_{A}=\mathbb{U}_{\Gamma(\mathcal{P})}(A)\otimes_{\mathbb{U}_{\Gamma(\mathcal{P})}(C_{\bullet})}\Omega_{\Gamma(\mathcal{P})}(C_{\bullet}).$
In particular, the $n$th Quillen homology module of $A$ is the $A$-module $\mathrm{HQ}_{n}(A)=\pi_{n}(\mathbf{L}_{A})$. The $n$th Quillen cohomology group of $A$ with coefficients in an $A$-module $M$ is the abelian group
$\mathrm{HQ}^{n}(A;M)=\pi^{n}\left(\operatorname{Hom}_{A\text{-}\mathrm{Mod}}(\mathbf{L}_{A},M)\right).$
@@
(1) Via the equivalences from Theorems 5.2 and 6.11, the statement follows from combining Proposition 7.6, Lemma 8.1, and Lemma 2.4.
(2) Via the equivalence from Theorem 6.11, the right adjoint $A\ltimes-$ is the forgetful functor
@@
The claim then follows from [Frankland15]*Proposition 3.40. Part (3) follows from part (1). ∎
@@ 9.   Example: The operad $\mathrm{Com}$
In this section, we apply our general constructions for Beck modules, universal enveloping algebra, Beck derivations and Kähler differentials to divided power algebras over the operad $\mathrm{Com}$ of associative, commutative (non-unital) algebras. We check that these correspond to the constructions obtained by the first author in [Dokas09, Dokas23].
@@ 9.1.   Classical definition and state of the art
Given an augmented $\mathbb{F}$-algebra $A$ with augmentation $\epsilon\colon A\to\mathbb{F}$, denote $A_{+}=\ker(\epsilon\colon A\to\mathbb{F})$.
@@
We follow the notation $A_{+}=\ker(\epsilon\colon A\to\mathbb{F})$ used in [Soublin87] and [Dokas09], though some authors use a different notation, notably $\overline{A}$ in [LodayV12]*§1.1.
@@
We define the ring $V(A)$ as the ring whose underlying $\mathbb{F}$-vector space is the tensor product $V(A)\coloneqq A\otimes_{\mathbb{F}}R_{f}$ and the multiplication is given by:
@@ 9.2.   Operadic point of view
Let us now use the characterisation of $\Gamma(\mathrm{Com})$-algebra from Theorem 3.3. Following [Ikonicoff20], if $B$ is a $\Gamma(\mathrm{Com})$-algebra, its multiplication is given by $\beta_{X_{2},(1,1)}$, and $\pi$, which represents the divided power operation $\gamma_{p}$, is given by $\beta_{X_{p},(p)}$.
@@
Note that $B$ injects into $\mathbb{U}_{\Gamma(\mathrm{Com})}(B)$, by identifying each $b\in B$ to $\beta_{X_{2},(1,1)}(b,-)$. For a $B$-module (i.e., a $\mathbb{U}_{\Gamma(\mathrm{Com})}(B)$-module) $M$, we will denote by $bm:=\beta_{X_{2},(1,1)}(b,m)=\beta_{X_{2},(1,1)}(b,-)\cdot m$, and $\pi(m):=\beta_{X_{p},(p)}(m)$, for $b\in B$, $m\in M$.
@@
By Proposition 7.2, a Beck derivation $d\colon B\to M$ is a linear map $d\colon B\to M$ satisfying $d(ab)=adb+bda$ and
@@
Since in characteristic $p$, $(p-1)!=-1$ (see Wilson’s Theorem [DummitF04]*§13.5 Exercise 6), one has:
@@
By Definition 7.3, the module of Kähler differentials $\Omega_{\Gamma(\mathrm{Com})}(B)$ of $B$ is the $\mathbb{U}_{\Gamma(\mathrm{Com})}(B)$-module generated by elements $db$ for $b\in B$, linear in $b$, under the relations:
$d(ab)$ $=d(\beta_{X_{2},(1,1)}(a,b))$
$=\beta_{X_{2},(1,1)}(-,b)\otimes da+\beta_{X_{2},(1,1)}(a,-)\otimes db$
$=bda+adb,$
@@
$\beta_{X_{p},(p-1,1)}(b,-)\otimes b$ $=\beta_{X_{2},(1,1)}(\beta_{X_{p-1},(p-1)}(b),-)\otimes db$
$=\frac{1}{(p-1)!}b^{p-1}db=-b^{p-1}db.$
@@
For all $i\in\{2,\dots,p-1\}$, we have again
$X_{p}=i\frac{1}{i}X_{p}=\sum_{\sigma\in\Sigma_{p-i}\times\Sigma_{i}/\Sigma_{p-i}\times\Sigma_{1}\times\Sigma_{i-1}}\frac{1}{i}X_{p}.$
So, $\beta_{X_{p},(p-i,i)}(b,-)=\sum_{\sigma\in\Sigma_{p-i}\times\Sigma_{i}/\Sigma_{p-i}\times\Sigma_{1}\times\Sigma_{i-1}}\frac{1}{i}X_{p}$ is one of the elements described in Definition 6.1 in point (7)
@@ 10.1.   Classical definition and state of the art
Following [Dokas04], denote by $w(L)$ the $\mathbb{F}$-vector space $R_{f}\otimes_{\mathbb{F}}u(L)$, where $R_{f}$ is the polynomial ring on one indeterminate $f$ as in Section 9. Then, $w(L)$ can be equipped with a ring structure such that $R_{f}\to w(L)$ and $u(L)\to w(L)$ are ring homomorphisms, and such that:
@@
$\operatorname{Der}_{p}(L^{\prime},(M,f))\coloneqq\{d\in\operatorname{Der}(L^{\prime},M):d(l^{[p]})=\underbrace{l\cdots l}_{p-1}dl+f(d(l)),\;l\in L^{\prime}\}$
@@
What we call the module of Kähler differentials of the Lie algebra $L$ is, by analogy with the case of commutative algebras, the $L$-Beck module $\Omega_{\operatorname{RLie}}(L)$ which represents the functor of Beck derivations. It follows from Theorem 10.8, Proposition 10.10 and [Pareigis68]*Lemma 2.1 that the module $\Omega_{\operatorname{RLie}}(L)$ of Kähler differentials is the $w(L)$-module $C(L)$ considered by Pareigis in [Pareigis68], cf. [Dokas04]*§1.3.
@@
The module of Kähler differentials of the restricted Lie algebra $L$, $\Omega_{\operatorname{RLie}}(L)$, is the $w(L)$-module with the following presentation: the generators are the symbols $dl$ for $l\in L$, and the relations are: